OpenStudy (90890908967777878888):

20 f+10 A.f(2+100 B.2(f+5) C.2(5+10)

OpenStudy (alexandervonhumboldt2):

20f+10 A: f*(2+100)=2f+100f B: 2*(f+5)=2f+10 C: 2*(5+10)=10+20=30 none of your choices equal 20f+10.

OpenStudy (90890908967777878888):

yes they r